# VidHal: Benchmarking Temporal Hallucinations in Vision LLMs

### v1

@@ Abstract
Vision Large Language Models (VLLMs) are widely acknowledged to be prone to hallucination. Existing research addressing this problem has primarily been confined to image inputs, with limited exploration of video-based hallucinations. Furthermore, current evaluation methods fail to capture nuanced errors in generated responses, which are often exacerbated by the rich spatiotemporal dynamics of videos. To address this, we introduce VidHal, a benchmark specially designed to evaluate video-based hallucinations in VLLMs. VidHal is constructed by bootstrapping video instances across common temporal aspects. A defining feature of our benchmark lies in the careful creation of captions which represent varying levels of hallucination associated with each video. To enable fine-grained evaluation, we propose a novel caption ordering task requiring VLLMs to rank captions by hallucinatory extent. We conduct extensive experiments on VidHal and comprehensively evaluate a broad selection of models. Our results uncover significant limitations in existing VLLMs regarding hallucination generation. Through our benchmark, we aim to inspire further research on I) holistic understanding of VLLM capabilities, particularly regarding hallucination, and II) extensive development of advanced VLLMs to alleviate this problem. The VidHal dataset and evaluation code are publicly available at https://github.com/Lookuz/VidHal.
@@ 1 Introduction
Building on the advancements of Large Language Models (LLMs), Vision LLMs (VLLMs) have recently gained significant attention. Models such as LLaVA [32, 33] have shown impressive performance across various visual understanding tasks involving both images and videos. Despite their potential, VLLMs are notably prone to hallucinations, where generated responses that appear to be plausible contradict the visual context [1]. This problem significantly compromises the reliability of VLLMs, hindering their practical use in real-world applications.
@@
To address this research gap, we present VidHal, a benchmark specifically designed to evaluate video-based hallucinations of VLLMs. VidHal features videos that comprehensively cover a wide range of temporal aspects, such as entity actions and sequence of events. Each video is automatically annotated with multiple captions exhibiting varying levels of aspect-specific hallucinations, capturing both subtle and significant discrepancies. In addition, we perform detailed human validation to ensure the robustness and reliability of our annotation process. An additional motivation stems from the limited metrics for quantifying hallucinations in VLLMs. To capture fine-grained hallucinatory errors of these models, we propose a unique caption ordering task that requires models to rank captions by hallucination levels. This consequently leads to a ranking-based NDCG metric and an MCQA accuracy metric, both are distinct from prior ones and specifically tailored to evaluate nuanced hallucinations in video-based VLLMs.
Using our VidHal dataset, we benchmark eleven VLLMs including both open-sourced and proprietary models, with abstracted results summarized in Figure 1. Through these extensive experiments, we identify limitations in nuanced video understanding among all evaluated VLLMs. Specifically, our findings reveal that existing VLLMs struggle to differentiate between captions with varying levels of hallucination. This deficiency is particularly evident when evaluating video-specific aspects, such as Direction and Order, as illustrated in Figure 1, indicating substantial room for improvement in current video-based VLLMs. Besides, we also notice that proprietary models, e.g., GPT-4o [41], frequently outperform open-source models by significant margins.
@@ 2.1 Vision Large Language Models
The emergence of powerful LLMs has advanced the development of VLLMs [32, 33, 27, 10, 56, 57, 55]. Typical methods in this category include LLaVA [32], mPLUG-Owl [57, 55, 56], InstructBLIP [10], and MiniGPT-4 [67]. These VLLMs rely on aligning vision encoders with LLMs using connective modules such as Q-Former [10, 26, 27, 61, 9] or MLPs [32, 33, 47] with the instruction tuning stage. Recent methods have extended visual inputs from images to (long) videos, delivering impressive joint spatial-temporal reasoning capabilities. For instance, VideoLLaMA2 [9] enhances the LLaMA model with video understanding capabilities through a Spatial-Temporal Convolution (STC) module. LLaVA-NeXT-Video [34, 62] presents an AnyRes approach that enables reasoning with long videos and high-resolution images.
@@ 2.2 Hallucinations in VLLMs
Despite their impressive performance on visual reasoning benchmarks, current VLLMs remain notoriously susceptible to hallucinations [20, 37, 68, 8]. A common demonstration is that the generated responses contain information which is inconsistent with the visual content [1, 35, 59, 53]. Most approaches address the hallucination problem with post-hoc techniques. For example, LURE [66] and Woodpecker [58] develop pipelines that assist VLLMs in revising their responses using expert models. To reduce bias from unimodal and statistical priors, contrastive decoding methods, such as VCD [24] and M3ID [13], along with attention calibration techniques like OPERA [18] are employed to refine token predictions. Building on the success of reinforcement learning for preference optimization [42], HA-DPO [63] and HIO [5] adopt this paradigm to fine-tune VLLMs, yielding outputs with fewer hallucinations. On the other hand, several benchmarks [29, 50, 48, 22, 30, 21, 51, 64, 8, 4, 52, 49, 7] have recently been constructed to quantify visual hallucinations, focusing primarily on object-based hallucinations in images.
@@ 2.3 Video Reasoning Benchmarks
The popularity of video-based VLLMs has been accompanied by the continual development of video benchmarks. Notable ones, such as SEEDBench [25], VideoBench [40], MVBench [28] and VideoMME [14], focus on instances featuring dynamic events that require temporal reasoning beyond a single frame. However, these benchmarks often suffer from a lack of diversity in both reasoning tasks and visual concepts. To address this, AutoEval-Video [6] and Perception Test [43] incorporate a greater variety of complex reasoning tasks, e.g., counterfactual and explanatory reasoning. TempCompass [38] enriches previous benchmarks with a wider range of temporal concepts. Notably, HallusionBench [17], FactVC [31] and VideoCon [2] partially cover video-based hallucinations. However, these benchmarks are limited by their focus on a few video-centric aspects and fail to capture nuanced hallucinations with their evaluation tasks. In contrast, our vidhal is specially designed to carefully evaluate these video-based hallucinations, encompassing a broad range of video aspects with tailored evaluation metrics.
@@ 3 VidHal Dataset Construction
We introduce VidHal, a video-language benchmark designed to evaluate hallucinations of Video-LLMs in a holistic manner. As depicted in Figure 2, VidHal comprises of video instances covering various temporal aspects to assess hallucinations across multiple dimensions and granularity of temporal reasoning.
@@ 3.1 Temporal Concept Selection
Hallucinations in image inputs can typically be classified into three types: object, attribute, and relation. In contrast, video-based hallucinations in VLLM responses may involve dynamic factors closely associated with the spatiotemporal patterns in video content. In light of this, we propose focusing on the following five aspects to ensure comprehensive coverage of the temporal concepts.
  • Attribute describes the fine-grained characteristics and properties of objects or subjects in the video. We additionally categorize this aspect into sub-aspects of Size, Shape, Colour, Count and State Change.
  • Object relates to the interactions between objects and entities within the video. We further delineate this aspect into two fine-grained sub-aspects: Object Recognition, identifying the objects engaged in interactions, and Interaction Classification, focusing on the semantics of how these objects interact.
  • Action refers to the movements and behaviours exhibited by entities in dynamic video scenes.
  • Direction indicates the orientation and movement trajectory of subjects or objects in the video.
  • Event Order represents the correct sequence of events in the video. During our collection, we retain videos that contain at least three distinct events.
We present one example illustrating the direction aspect in Figure 2, with additional examples available in the supplementary material.
@@ 3.2 Hallucinatory Caption Generation
Based on the aspects in Section 3.1, we build our benchmark upon four public video understanding datasets: TempCompass [38], Perception Test [43], MVBench [28] and AutoEval-Video [6]. TempCompass and MVBench provide extensive coverage across all five temporal aspects. Perception Test includes a broader selection of human-object interactions, and AutoEval-Video emphasizes scenarios with attribute changes.
Existing hallucination benchmarks [29, 50] predominantly rely on binary questions for evaluation, which limits their effectiveness for assessing hallucinations in video-based VLLMs. Unlike static images, video content features complex spatiotemporal dynamics that can lead to subtle hallucinations in VLLMs, e.g., minor inconsistencies in events or motion. These nuanced errors, however, are difficult to detect through binary questions.
To address this issue, we advocate a novel evaluation protocol that incorporates several carefully annotated captions. Specifically, each video will be annotated with a set of $M$ captions that reflect varying degrees of hallucination in VLLMs. Given the cost and labor intensity of manual annotation, we opt for automatic caption generation using a carefully designed pipeline illustrated in Figure 2.
Anchor Caption Generation. The video instances in VidHal are sourced from various public datasets, resulting in distinct associated metadata such as long-form captions in AutoEval-Video and question-answer pairs in MVBench and TempCompass. To ensure structure consistency and information granularity in the respective dataset description across all instances, we automatically generate an anchor caption for each video. Specifically, we input the metadata for each video $V^{i}$ into GPT-4o and prompt it to generate a concise and accurate description $y^{i}_{+}$ using the provided metadata information.
@@ 3.3 Dataset Statistics and Human Validation
FactVC [31] Video Captioning FactVC
VideoCon [2] VLE, NLE ROC-AUC, $Q^{2}$ score
@@
VidHal(Ours) MCQA, Caption Ordering Accuracy, NDCG
Table 1: Comparison of our benchmark dataset with typical video-based datasets. VLE and NLE represent the tasks of visual-language entailment and natural language explanation, respectively.
@@ 4.1 Evaluation Tasks
Notably, our relative caption ordering task is more challenging than previous binary questions. This complexity arises from certain paired questions in our dataset where both options contain hallucinatory information, making it more difficult to distinguish them apart as opposed to ⟨positive, hallucinatory⟩ pairs.
@@ 4.2 Evaluation Metrics
For the MCQA setting, we adopt the standard accuracy metric,
@@
Inspired by metrics from the information retrieval domain [15], we adapt the well-established Normalized Discounted Cumulative Gain (NDCG) [19] for hallucination assessment in VidHal. Unlike previous metrics like POPE [29], our metric awards partial credit for correctly ordered caption pairs even when the optimal ranking is not achieved. As such, we expect the metric to effectively capture both subtle and severe hallucinations generated by video-based VLLMs. Formally, we define our adapted NDCG metric as follows:
@@
where $\hat{y}^{i,j}$ represents $j^{th}$ caption in the ranked order predicted by the VLLM. The perfect ordering is achieved when $\hat{y}^{i,1}=y^{i}_{+}$ and $\{\hat{y}^{i,j}=y^{i,j-1}_{-}\}_{j=2\to M}$. To evaluate predicted caption orders relative to this ideal sequence, a relevance function $r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)$ is designed to assign higher scores to $\hat{y}^{i,j}$ with lower hallucinatory extent. We formalize the relevance function below:
@@
where $\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})$ denotes the position of $\hat{y}^{i,j}$ in $\mathcal{Y}^{i}_{*}$. Finally, $\text{DCG}_{i}$ is normalized to a range of 0 to 1 by introducing $\text{iDCG}_{i}$ and $\text{rDCG}_{i}$. These two terms represent the maximum and minimum $\text{DCG}_{i}$ scores obtained from the optimal ordering $\mathcal{Y}^{i}_{*}$ and its reverse, respectively,

### v2

@@ Abstract
Vision Large Language Models (VLLMs) are widely acknowledged to be prone to hallucinations. Existing research addressing this problem has primarily been confined to image inputs, with limited exploration of video-based hallucinations. Furthermore, current evaluation methods fail to capture nuanced errors in generated responses, which are often exacerbated by the rich spatiotemporal dynamics of videos. To address this, we introduce VidHal, a benchmark specially designed to evaluate video-based hallucinations in VLLMs. VidHal is constructed by bootstrapping video instances across a wide range of common temporal aspects. A defining feature of our benchmark lies in the careful creation of captions which represent varying levels of hallucination associated with each video. To enable fine-grained evaluation, we propose a novel caption ordering task requiring VLLMs to rank captions by hallucinatory extent. We conduct extensive experiments on VidHal and comprehensively evaluate a broad selection of models. Our results uncover significant limitations in existing VLLMs regarding hallucination generation. Through our benchmark, we aim to inspire further research on I) holistic understanding of VLLM capabilities, particularly regarding hallucination, and II) extensive development of advanced VLLMs to alleviate this problem. Our VidHal dataset and evaluation code are publicly available at https://github.com/Lookuz/VidHal.
@@ 1 Introduction
Building on the advancements of Large Language Models (LLMs), Vision LLMs (VLLMs) have recently gained significant attention. Models such as LLaVA [33, 34] have shown impressive performance across various visual understanding tasks involving both images and videos. Despite their potential, VLLMs are notably prone to hallucinations, where generated responses that appear to be plausible contradict the visual context [1, 57]. This problem significantly compromises the reliability of VLLMs, hindering their practical use in real-world applications.
@@
To address this research gap, we present VidHal, a benchmark specifically designed to evaluate video-based hallucinations of VLLMs. VidHal features videos that comprehensively cover a broad range of temporal aspects, such as entity actions and sequence of events. Each video is automatically annotated with multiple captions exhibiting varying levels of aspect-specific hallucinations, capturing both subtle and significant discrepancies. In addition, we perform detailed human validation to ensure the robustness and reliability of our annotation process. An additional motivation stems from the limited metrics for quantifying hallucinations in VLLMs. To capture fine-grained hallucinatory errors of these models, we propose a unique caption ordering task that requires models to rank captions by hallucination levels. This consequently leads to a ranking-based NDCG metric and an MCQA accuracy metric, both are distinct from prior ones and specifically tailored to evaluate nuanced hallucinations in video-based VLLMs.
Using our VidHal dataset, we benchmark thirteen VLLMs including both open-sourced and proprietary models, with abstracted results summarized in Figure 1. Through these extensive experiments, we identify limitations in nuanced video understanding among all evaluated VLLMs. Specifically, our findings reveal that existing VLLMs struggle to differentiate between captions with varying levels of hallucination. This deficiency is particularly evident when evaluating video-specific aspects, such as Direction and Order, as illustrated in Figure 1, indicating substantial room for improvement in current video-based VLLMs. Besides, we also notice that proprietary models, e.g., GPT-4o [42], frequently outperform open-source models by significant margins.
@@ 2.1 Vision Large Language Models
The emergence of powerful LLMs has advanced the development of VLLMs [33, 34, 26, 9, 60, 61, 59]. Typical methods in this category include LLaVA [33], mPLUG-Owl [61, 59, 60], InstructBLIP [9], and MiniGPT-4 [73]. These VLLMs rely on aligning vision encoders with LLMs using connective modules such as Q-Former [9, 25, 26, 65, 8] or MLPs [33, 34, 48] with the instruction tuning stage. Recent methods have extended visual inputs from images to (long) videos, delivering impressive joint spatial-temporal reasoning capabilities. For instance, VideoLLaMA2 [8] enhances the LLaMA model with video understanding capabilities through a Spatial-Temporal Convolution (STC) module. LLaVA-NeXT-Video [35, 66] presents an AnyRes approach that enables reasoning with long videos.
@@ 2.2 Hallucinations in VLLMs
Despite their impressive performance on visual reasoning benchmarks, current VLLMs remain notoriously susceptible to hallucinations [19, 38, 74, 7]. A common demonstration is that the generated responses contain information which is inconsistent with the visual content [1, 36, 63, 55]. Most approaches address the hallucination problem with post-hoc techniques. For example, LURE [71] and Woodpecker [62] develop pipelines that assist VLLMs in revising their responses using expert models. To reduce bias from unimodal and statistical priors, contrastive decoding methods, such as VCD [23] and M3ID [12], along with attention calibration techniques like OPERA [17] are employed to refine token predictions. Building on the success of reinforcement learning for preference optimization in LLM development [43], HA-DPO [67], POVID [70] and CSR [72] adopt this paradigm to fine-tune VLLMs, yielding outputs with fewer hallucinations.
@@ 2.3 Video Reasoning Benchmarks
The rise of video-based VLLMs has driven the development of numerous video benchmarks. Notable examples, such as SEEDBench [24], VideoBench [41], MVBench [28], and VideoMME [13], focus on dynamic events requiring temporal reasoning beyond individual frames. However, these benchmarks often lack diversity in reasoning tasks and visual concepts. To address this, AutoEval-Video [5] and Perception Test [44] introduce complex reasoning tasks such as counterfactual and explanatory reasoning, while TempCompass [39] expands temporal concept coverage. Several benchmarks [29, 51, 49, 21, 31, 20, 53, 68, 7, 4, 54, 50, 6] have been constructed to quantify visual hallucinations, primarily targeting object-based hallucinations in images. HallusionBench [16], VideoCon [2], and Vript [58] provides partial coverage of video-based hallucinations, while VideoHallucer [52] introduces a benchmark for hallucination detection in videos. However, these benchmarks provide limited coverage of spatio-temporal concepts, focusing on conventional aspects like actions and objects while neglecting other video-centric elements such as direction. Additionally, their evaluation strategies primarily follow image-based approaches, which we argue are less effective in capturing nuanced, video-specific hallucinations.
@@ 3 VidHal Dataset Construction
We introduce VidHal, a unique video-language benchmark designed to evaluate hallucinations of Video-LLMs in a comprehensive manner. As depicted in Figure 2, VidHal comprises of video instances which span a diverse spectrum of temporal aspects, including previously unexplored aspects such as directional movement. In contrast to previous studies on video hallucination evaluation [58, 52, 2], VidHal incorporates multiple hallucinated captions per video, enabling the assessment of video hallucinations at multiple levels of granularity.
3.1 Temporal Hallucinations in Videos
Hallucinations in VLLMs occur when the model fabricates details in its responses that contradict the provided visual content. Compared to images, video hallucinations extend beyond static visual elements to include misperceptions of dynamic changes within scenes. We categorize these temporal hallucinations into two semantic levels:
Lexical Semantics (L-Sem) captures instances where VLLMs misinterpret words related to temporal features, including nouns referring to objects or attributes (e.g., misidentifying a color change from green to red as green to orange) and verbs describing actions (e.g., interpreting ”kicking a ball” as ”throwing a ball”).
Clause Semantics (C-Sem) encompasses errors involving event descriptions and their sequences, where the VLLM incorrectly predicts the order of events occurring in the video. For example, given sequentially occurring events $A$ and $B$ in a video, the model may perceive $B$ preceding $A$.
By addressing these two dimensions of video-based hallucinations, VidHal offers holistic coverage over the level of detail in which VLLMs may hallucinate.
@@ 3.2 Temporal Concept Selection
Prior research on hallucination evaluation for both images [29, 51, 46] and videos [52, 58, 16] has predominantly focused on common visual aspects such as action- and object-based hallucinations. However, video-based hallucinations may involve additional dynamic factors associated with spatio-temporal patterns, which are overlooked in these studies. In light of this, we propose to focus on the following five aspects to ensure comprehensive coverage of temporal concepts. Specifically, the first four aspects address hallucinations based on lexical semantics, while the fifth targets those based on clause-level semantics.
  • Attribute (L-Sem) describes the fine-grained characteristics and properties of objects or subjects in the video. We additionally categorize this aspect into sub-aspects of Size, Shape, Color, Count and State Change.
  • Object (L-Sem) relates to the interactions between objects and entities within the video. We further delineate this aspect into two fine-grained sub-aspects: Object Recognition, identifying the objects engaged in interactions, and Interaction Classification which concentrate on how these objects interact with other objects or subjects.
  • Action (L-Sem) refers to the movements and behaviours exhibited by entities in dynamic video scenes.
  • Direction (L-Sem) indicates the orientation and movement trajectory of subjects or objects in the video.
  • Event Order (C-Sem) represents the correct sequence of events in the video. During our collection, we retain videos that contain at least three distinct events.
We present an example that illustrates the direction aspect in Figure 2, with additional examples available in the supplementary material.
@@ 3.3 Hallucinatory Caption Generation
Based on the aspects in Section 3.2, we build our benchmark upon four public video understanding datasets: TempCompass [39], Perception Test [44], MVBench [28] and AutoEval-Video [5]. TempCompass and MVBench extensively cover all five temporal aspects, while Perception Test and AutoEval-Video highlights human-object interactions and attribute changes, respectively.
Existing hallucination benchmarks [29, 51] rely mostly on binary questions for evaluation, limiting their efficacy in detecting subtle video hallucinations, such as minor event inconsistencies. To address this issue, we advocate a novel evaluation protocol incorporating several carefully annotated captions. Specifically, each video will be annotated with a set of $M$ captions that reflect varying degrees of hallucination in VLLMs. Given the cost and labor intensity of manual annotation, we opt for automatic caption generation using a carefully designed pipeline illustrated in Figure 2.
Anchor Caption Generation. The video instances in VidHal are sourced from various public datasets, resulting in distinct associated metadata such as long-form captions in AutoEval-Video and question-answer pairs in MVBench. To ensure structure consistency and information granularity in the respective dataset description across all instances, we automatically generate an anchor caption for each video. Specifically, we input the metadata for each video $V^{i}$ into GPT-4o and prompt it to generate a concise and accurate description $y^{i}_{+}$ using the provided metadata information.
@@ 3.4 Dataset Statistics and Human Validation
Size Shape Color Count State-Change Recognition Interaction
@@
Vript [58] Video Captioning F1 Score
Event Ordering Accuracy
VideoCon [2] VL Entailment ROC-AUC
@@
VidHal(Ours) MCQA Accuracy
Caption Ordering NDCG
Table 1: Comparison of our benchmark dataset with existing video-based reasoning and hallucination evaluation datasets. For datasets with multiple evaluation tasks, only those relevant to hallucination evaluation are included. VL Entailment denotes the task of video-language entailment, while Event Ordering prompts the model to determine the chronological sequence of scenes in a video.
@@ 4.1 Evaluation Tasks
Notably, our relative caption ordering task is more challenging than previous binary questions. This complexity arises from certain paired questions in VidHal where both options are hallucinatory, making them harder to distinguish as opposed to ⟨positive, hallucinatory⟩ pairs.
Model Release Date Vision Encoder LLM #Params #Frames Accuracy NDCG
Naive Relative
Baseline
Random - - - - - 0.326 0.505 0.480
Open-Sourced Models
VideoChat [27] May 2023 EVA-CLIP-G Vicuna 7B 8 0.381 0.475 0.488
LLaMA-VID [30] Nov. 2023 EVA-CLIP-G Vicuna 7B 1fps 0.358 0.486 0.521
VideoChat2 (Vicuna) [28] Nov. 2023 UMT-L Vicuna 7B 16 0.426 0.486 0.577
VideoChat2 (Mistral) May 2024 UMT-L Mistral 7B 16 0.443 0.503 0.475
VideoChat2 (Phi) Jun. 2024 UMT-L Phi3 3.8B 16 0.514 0.626 0.612
mPLUG-Owl3 [59] Aug. 2024 SigLIP/SO400M Qwen2 7B 16 0.596 0.641 0.707
LLaVA-NeXT-Video (7B) [66] Apr. 2024 SigLIP/SO400M Vicuna 7B 32 0.509 0.518 0.620
LLaVA-NeXT-Video (32B) Apr. 2024 SigLIP/SO400M Qwen1.5 32B 32 0.663 0.641 0.747
VideoLLaMA2 (7B) [8] Jun. 2024 CLIP ViT-L/14 Mistral 7B 8 0.541 0.564 0.622
VideoLLaMA2 (72B) Jun. 2024 CLIP ViT-L/14 Qwen2 72B 8 0.647 0.787 0.760
Proprietary Models
GPT-4o [42] May 2024 - - - 1fps 0.772 0.840 0.826
Gemini-1.5 (Flash) [45] May 2024 - - - 1fps 0.657 0.738 0.745
Gemini-1.5 (Pro) May 2024 - - - 1fps 0.671 0.765 0.753
@@ 4.2 Evaluation Metrics
We employ the standard accuracy metric:
@@
Inspired by metrics from the information retrieval domain [14], we adapt the well-established Normalized Discounted Cumulative Gain (NDCG) [18] for hallucination assessment in VidHal. Unlike previous metrics like POPE [29], our metric awards partial credit for correctly ordered caption pairs even when the optimal ranking is not achieved. As such, we expect the metric to effectively capture and distinguish both subtle and severe hallucinations generated by video-based VLLMs. Formally, we define our adapted NDCG metric as follows:
@@
and $\hat{y}^{i,j}$ represents $j^{th}$ caption in the ranked order predicted by the VLLM. The perfect ordering is achieved when $\hat{y}^{i,1}=y^{i}_{+}$ and $\{\hat{y}^{i,j}=y^{i,j-1}_{-}\}_{j=2\to M}$. To evaluate predicted caption orders relative to this ideal sequence, a relevance function $r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)$ is designed to assign higher scores to $\hat{y}^{i,j}$ with lower hallucinatory extent.
@@
where $\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})$ denotes the position of $\hat{y}^{i,j}$ in $\mathcal{Y}^{i}_{*}$. Finally, $\text{DCG}_{i}$ is normalized to a range of $[0,1]$ using $\text{iDCG}_{i}$ and $\text{rDCG}_{i}$, with a score of 1 indicating perfect alignment of the predicted order with $\mathcal{Y}^{i}_{*}$. Specifically, these terms represent the maximum and minimum $\text{DCG}_{i}$ scores obtained from the optimal ordering $\mathcal{Y}^{i}_{*}$ and its reverse, respectively,